OpenStudy (anonymous):

simplify each expression . justify your steps ` 1. b + 2 + 5b 2. 16 + 5b + 3b + 9 3. 5a + 6t + 9 + 2a

OpenStudy (anonymous):

1. 6b+2 2.8b+25 3.7a+6t+9

OpenStudy (anonymous):

1. 6b+2= 2(3b+1)

OpenStudy (anonymous):

For each one we try and combine like variables

OpenStudy (anonymous):

2. 8b+25 3. 7a+6t+9

OpenStudy (anonymous):

just add them all the same variables add them together. like exmple b+5b=6b :]